10) Use 8-bit representations to compute 57 + (-118)
Answer #1

hello,

let us solve this question.

57 in 8 bit binary = 00111001

118 in 8 bit binary=01110110, so taking 2's complement of 118 will give binary of (-118)

so 2's complement of (01110110)=(10001001)+(00000001)=(10001010)   i.e, binary of (-118)

now we can add binary of 57 and (-118):

(00111001)+(10001010)=(11000011)     the left most bit in the answer is 1 which indicates it as a negative value , so we will take 2's complement of answer and the final answer will be -ve of 2's complement of (11000011)

2's complement of (11000011)=(00111100)+(00000001)=(00111101) which is 61 in decimal.

So the final answer is (-61)

i hope i was able to solve your problem to a greater extent, please feel free to comment your queries, Please consider my efforts and upvote my solution.

Thanku:)